# Toward a Multi-Echelon Cyber Warfare Theory: A Meta-Game-Theoretic Paradigm for Defense and Dominance

### v1

@@ 3.2 Role of Game Theory in Cyber Warfare
Risks in cyber warfare arise from adversarial interactions, human and organizational weaknesses, systemic dependencies, and vulnerabilities at multiple levels of strategy. Game theory offers a rigorous framework to model these risks by analyzing the strategic interplay between attackers, defenders, and third parties within broader environments yang2024prada . In the following, we outline key categories of risks and their game-theoretic implications.
@@
Attackers exploit vulnerabilities and disrupt systems, while defenders deploy countermeasures to contain threats. However, defensive actions carry their own risks: poorly designed responses can escalate conflict, produce unintended consequences, or cause collateral damage to civilian infrastructure or neutral actors. Game-theoretic models capture these dynamics as contests of adaptation. For example, Stackelberg security games illustrate how commitment to patrols or monitoring schedules can deter adversaries, but overly aggressive commitments can backfire if attackers adapt. Similarly, signaling games explain how misinterpreted defensive signals can escalate tensions or induce attackers to test hidden capabilities.
@@
Weaknesses within organizations introduce exploitable openings. Human errors, such as misconfigurations, delayed responses, or insider threats, intersect with resource constraints such as underfunded budgets or inadequate training. Failures of coordination between technical, legal, and executive teams compound risks. Game theory highlights how internal misalignments shape external vulnerabilities. Principal–agent models capture misaligned incentives within organizations (e.g., between management and frontline responders), while Bayesian games model how uncertainty about insider intentions leads to costly monitoring and screening.
@@
Risks manifest differently across the echelons. At the technical level, software bugs or unpatched systems create immediate vulnerabilities. At the tactical level, operational errors, such as mistargeting or premature exposure, undermine missions. Operational-level failures include misaligned objectives or communication overload. At the strategic level, overextension and misjudgment of adversaries compromise campaigns. At the grand strategic level, long-term risks include sanctions, political fallout, or escalation to kinetic conflict. A hierarchical game framework illustrates how local actions (e.g., disabling a power grid) propagate upward into operational disruptions, strategic instability, and national security repercussions. For instance, dynamic stochastic games can model how a tactical intrusion cascades into operational degradation, while differential games highlight escalation dynamics over extended campaigns.
@@
Global interdependence magnifies risks. Third-party software and hardware introduce supply chain vulnerabilities, while geopolitical uncertainty and the behavior of neutral actors complicate planning. Reliance on shared infrastructures such as DNS systems or cloud providers creates systemic exposure, where a single failure cascades across multiple stakeholders. Game theory models these as multi-player games with externalities, where equilibrium outcomes are inefficient unless actors coordinate. Public goods games capture underinvestment in shared security (e.g., collective defense of Internet routing), while network games explain how localized disruptions propagate through interdependent systems.
@@
Game theory enables organizations to quantify risks, anticipate adversarial moves, balance trade-offs, and improve decision-making under uncertainty. Payoff structures capture expected costs and benefits, guiding the choice of dominant or mixed strategies that mitigate risks while preserving effectiveness. Importantly, game theory also provides metrics, such as expected loss, value of the game, or the window of superiority raio2023toward , that quantify the consequences of different strategies and the efficiency of equilibrium outcomes. For example, Parrondo’s paradox shows how the alternation of weak defenses individually can produce a strong overall strategy, while Braess’ paradox warns that adding defensive resources can inadvertently worsen systemic performance.
Anticipation and adaptation are central: honeypot games demonstrate how deception disrupts attacker reasoning, while adaptive security games model defenders that evolve in response to emerging threats. Trade-off analysis directs resource allocation, for instance between detection tools and endpoint protection, under budget constraints. Finally, systematic modeling reduces the reliance on intuition when adversary capabilities or intentions remain uncertain and highlights the value of coordination across sectors to pool resources and strengthen collective defense.
@@
Clausewitz underscored this distinction: “War is not the action of a living force upon a lifeless mass ... but always the collision of two living forces” clausewitz1976onwar . Unlike mechanical systems with predictable responses, warfare unfolds as a strategic interaction in which each move is met by a countermove, generating complexity, unpredictability, and frequent reversals.
@@
Two losing games, such as a biased coin toss (Game A) and a capital-dependent coin toss (Game B), when alternated or randomly mixed, can form a winning strategy. This counterintuitive effect highlights how deception strategies that appear weak in isolation may, when combined, generate advantage over adaptive adversaries. In harmer1999losing , the authors demonstrate how combining two individually losing strategies can paradoxically yield a winning outcome. Translated into cyber defense, this suggests that alternating between deceptive tactics, each of which may appear ineffective or even detrimental in isolation, can produce an overall advantage by disrupting attacker expectations and decision-making. For instance, switching between high-visibility decoys and low-profile honeypots may confuse adversaries, forcing them into suboptimal choices. This insight highlights the value of counterintuitive strategy design in cyber warfare: by embracing paradox and leveraging surprise, defenders can transform apparent weaknesses into systemic strengths.
@@
More information is not always beneficial. Akerlof’s “Market for Lemons” demonstrates that full transparency can collapse markets akerlof1978market , while Gossner’s work shows that partial ignorance can enable trade gossner2000comparison . In cyber warfare, defenders who release too much information about their posture may strengthen attackers, while carefully designed opacity can restore balance.
@@
In network routing, adding capacity can worsen overall performance at equilibrium. In a road network, removing a link improved travel times for all drivers di2014braess . Similarly, in the cyber battlefield, deploying additional defensive layers or controls in networks can sometimes degrade resilience by creating new bottlenecks or redirecting attacks.
@@
Surprise is central to cyber warfare, shaping strategy and amplifying operational impact. It exploits adversarial blind spots and introduces unexpected disruptions that undermine defenses. Unlike conventional warfare, where surprise is achieved through maneuvers or speed, cyber operations leverage stealth, misinformation, and novel attack vectors. Advanced Persistent Threats (APTs) exemplify this, infiltrating quietly and striking only when conditions maximize the effect. Zero-day exploits deliver shock similarly by exploiting unknown vulnerabilities before defenses exist.
@@
Game-theoretic approaches formalize this interplay of deception and surprise. Cyber deception games, for instance, model how defenders mislead attackers about intentions and capabilities. In yang2025bilevel , cognitive triggers are designed to create microdeceptions that exploit bounded rationality, inducing adversaries into suboptimal choices by shaping their beliefs and expectations. This slows the attack process, forces attackers to waste additional effort, and ultimately creates strategic advantages for the defender.
Figure 1: A multi-echelon cyber warfare framework that outlines the hierarchical structure of cyber conflict, moving from tactical-level actions to cyber policy-level strategies. It provides a staged progression of cyber operations, illustrating how China’s cyber warfare campaign against Taiwan would unfold across different strategic layers.
@@ 4 Multi-Echelon Game-Theoretic Paradigm
Game theory provides the mathematical rigor to represent this hierarchy. Each echelon, policy, strategic, operational, tactical, and technical, corresponds to a different game with its own players, actions, payoffs, and equilibrium concepts. Yet, these games are not isolated: Equilibria at one level act as constraints or objectives for the next, while outcomes at lower levels feed back to reshape higher-level decisions. This recursive dependence gives rise to what we term a meta-game, in which stability requires consistency across all levels.
@@ 4.1 Multi-Echelon Games
At the top of the hierarchy are cyber policy decisions. These include international agreements (e.g., cyber alliances, norms of behavior, and frameworks like the Tallinn Manual) and investments in national infrastructure. Such interactions can be modeled with coalition games aumann1974cooperative . In this setting, the players are countries, $N=\{1,\dots,n\}$. Any subgroup of countries $S\subseteq N$ can form a coalition, and the coalition value function $v(S)$ describes the benefit of cooperation (e.g., shared defense, intelligence sharing). Game-theoretic solution concepts such as the core or the Shapley value tell us how to fairly divide the benefits of cooperation so that alliances remain stable.
@@
At the strategic level, cyber campaigns unfold over multiple stages, indexed by
$k\in\{k_{1},k_{2},\dots,K\},$
as shown in Figure 1. Each stage corresponds to a different phase of operations: reconnaissance ($k=k_{1}$), disruption ($k=k_{2}$), escalation ($k=k_{3}$), sustained assault ($k=K$), and finally strategic messaging.
@@
As shown in the figure, the sequences of actions over time determine whether escalation occurs at stage $k=k_{3}$. Here, equilibrium concepts such as the $\epsilon$-Perfect Bayesian Nash Equilibrium huang2020dynamic are important: they model how attackers form beliefs when they are uncertain about which deception mechanism $\lambda$ is active.
@@
Finally, at the technical level, the focus is on developing new technological and computational solutions that directly enhance cyber defense. Large Language Models (LLMs) and agentic AI serve as key enablers, providing the ability to generate adaptive responses, automate decision support, and synthesize vast streams of threat intelligence. For instance, LLM-driven agents can monitor logs, correlate signals across domains, and propose countermeasures in real time, while multi-agent AI frameworks enable distributed coordination among defensive components zhu2025gamellm . In addition, adversarial learning techniques, such as Generative Adversarial Networks (GANs), support the creation of realistic honeypots and synthetic data sets that improve detection and resilience gabrys2024honeygan . At this level, the emphasis is on computational innovation that empowers defenders with tools capable of anticipating, deceiving, and outmaneuvering adversaries in highly dynamic environments.
@@ 4.2 Cross-Echelon Relationships
A canonical example of this setup is the Colonel Blotto game roberson2006colonel ; hart2008discrete , in which two players simultaneously distribute forces across multiple battlefields, and outcomes at each battlefield are determined by relative strength. Here, the “battlefields” correspond to operational categories $\Lambda$, and the policy-derived weights $w^{*}$ adjust the relative importance of each category. Equilibrium allocations determine which operations dominate, are neutralized, or remain contested.
@@
At the operational level, a campaign is broken into phases $k=1,\dots,K$, each with a state $s^{k}$. The attacker chooses actions $a_{a}^{k}$ with strategy $\pi_{a}^{k}(s^{k})$, while the defender chooses actions $a_{d}^{k}$ with strategy $\pi_{d}^{k}(s^{k},\lambda)$, where $\lambda$ indexes which deception mechanism (e.g., honeypot, decoy) is deployed gabrys2024honeygan ; huang2019adaptive ; huang2020farsighted ; yang2025timing ; yang2025timing . The defender’s cumulative payoff is
@@
At the technical level, $\theta$ captures the state of enabling technologies (e.g., advanced LLMs or GAN-based deception systems). Improvements at this level reduce costs $c_{d}(\theta)$, expand feasible action sets $A_{d}(\theta)$, alter the equilibrium strategies $(\xi_{d}^{*}(\theta),\xi_{a}^{*}(\theta))$ and their payoffs $U_{d}^{T},U_{a}^{T}$, and change the transition dynamics $T(\cdot;\theta)$. The associated equilibrium is denoted $E_{\text{technical}}=\theta^{*},$ representing the dominant technological state that parameterizes the equilibria at all higher levels. In this way, advances at the technical level propagate upward, shifting equilibria across every higher echelon of the hierarchy.

### v2

@@ 3.2 Role of Game Theory in Cyber Warfare
Risks in cyber warfare arise from adversarial interactions, human and organizational weaknesses, systemic dependencies, and vulnerabilities at multiple levels of strategy. Game theory provides a rigorous framework for modeling these risks by analyzing the strategic interplay among attackers, defenders, and third parties within broader environments Yang et al. (2025c). In the following, we outline key categories of risks and their game-theoretic implications.
@@
Attackers exploit vulnerabilities and disrupt systems, while defenders deploy countermeasures to contain threats. However, defensive actions carry their own risks: poorly designed responses can escalate conflict, produce unintended consequences, or cause collateral damage to civilian infrastructure or neutral actors. Game-theoretic models capture these dynamics as contests of adaptation. For example, Stackelberg security games, An et al. (2017); Yang and Zhu (2025b) illustrate how commitment to patrols or monitoring schedules can deter adversaries, but overly aggressive commitments can backfire if attackers adapt. Similarly, signaling games Li and Zhu (2023); Pawlick et al. (2018b) explain how misinterpreted defensive signals can escalate tensions or induce attackers to test hidden capabilities.
@@
Weaknesses within organizations introduce exploitable openings. Human errors, such as misconfigurations, delayed responses, or insider threats, intersect with resource constraints such as underfunded budgets or inadequate training. Coordination failures among technical, legal, and executive teams compound risks. Game theory highlights how internal misalignment shapes external vulnerabilities. Principal-agent models Zhang and Zhu (2021b); Chen and Zhu (2018) capture misaligned incentives within organizations (e.g., between management and frontline responders), while Bayesian games model how uncertainty about insider intentions leads to costly monitoring and screening Huang and Zhu (2021b, 2019b).
@@
Risks manifest differently across the echelons. At the technical level, software bugs or unpatched systems create immediate vulnerabilities. At the tactical level, operational errors, such as mistargeting or premature exposure, undermine missions. Operational-level failures include misaligned objectives or communication overload. At the strategic level, overextension and misjudgment of adversaries compromise campaigns. At the grand strategic level, long-term risks include sanctions, political fallout, and the risk of escalation to kinetic conflict. A hierarchical game framework illustrates how local actions (e.g., disabling a power grid) propagate upward into operational disruptions, strategic instability, and national security repercussions. For instance, dynamic stochastic games can model how a tactical intrusion cascades into operational degradation Huang and Zhu (2021a); Zhu and Başar (2009), while differential games highlight escalation dynamics over extended campaigns Huang et al. (2020).
@@
Global interdependence magnifies risks. Third-party software and hardware introduce supply chain vulnerabilities, while geopolitical uncertainty and the behavior of neutral actors complicate planning. Reliance on shared infrastructures such as DNS systems or cloud providers creates systemic exposure, where a single failure cascades across multiple stakeholders. Game theory models these as multi-player games with externalities, where equilibrium outcomes are inefficient unless actors coordinate. Public goods games Kosfeld et al. (2009); Galbiati and Vertova (2008) capture underinvestment in shared security (e.g., collective defense of Internet routing), while network games explain how localized disruptions propagate through interdependent systems Liu et al. (2023, 2022).
@@
Game theory enables organizations to quantify risks, anticipate adversarial moves, balance trade-offs, and improve decision-making under uncertainty. Payoff structures capture expected costs and benefits, guiding the choice of dominant or mixed strategies that mitigate risks while preserving effectiveness. Importantly, game theory also provides metrics, such as expected loss, value of the game, or the window of superiority Raio (2023), that quantify the consequences of different strategies and the efficiency of equilibrium outcomes. For example, Parrondo’s paradox shows how the alternation of weak defenses individually can produce a strong overall strategy Harmer and Abbott (2002), while Braess’ paradox warns that adding defensive resources can inadvertently worsen systemic performance Nagurney and Nagurney (2021).
Anticipation and adaptation are central: honeypot games demonstrate how deception disrupts attacker reasoning, while adaptive security games model defenders that evolve in response to emerging threats Huang and Zhu (2020); Zhu et al. (2025); Píbil et al. (2012). Trade-off analysis directs resource allocation, for instance between detection tools and endpoint protection, under budget constraints. Finally, systematic modeling reduces the reliance on intuition when adversary capabilities or intentions remain uncertain and highlights the value of coordination across sectors to pool resources and strengthen collective defense.
@@
Clausewitz stated this distinction: “War is not the action of a living force upon a lifeless mass ... but always the collision of two living forces” von Clausewitz (1976). Unlike mechanical systems with predictable responses, warfare unfolds as a strategic interaction in which each move is met by a countermove, generating complexity, unpredictability, and frequent reversals.
@@
Two losing games, such as a biased coin toss (Game A) and a capital-dependent coin toss (Game B), when alternated or randomly mixed, can form a winning strategy. This counterintuitive effect highlights how deception strategies that appear weak in isolation may, when combined, generate advantage over adaptive adversaries. In Harmer and Abbott (1999), the authors demonstrate how combining two individually losing strategies can paradoxically yield a winning outcome. Translated into cyber defense, this suggests that alternating between deceptive tactics, each of which may appear ineffective or even detrimental in isolation, can produce an overall advantage by disrupting attacker expectations and decision-making. For instance, switching between high-visibility decoys and low-profile honeypots may confuse adversaries, forcing them into suboptimal choices. This insight highlights the value of counterintuitive strategy design in cyber warfare: by embracing paradox and leveraging surprise, defenders can transform apparent weaknesses into systemic strengths.
@@
More information is not always beneficial. Akerlof’s “Market for Lemons” shows how asymmetric information can cause market collapse even when greater transparency is introduced Akerlof (1978). Gossner demonstrates that different information structures can strictly change equilibrium outcomes and that partial ignorance may, in some environments, sustain trade or improve incentives Gossner (2000). These results illustrate a deeper structural insight: increasing information does not monotonically improve performance; instead, it reshapes strategic behavior.
In cyber warfare, this paradox emerges in disclosure strategy. Defenders who reveal too much about their defensive posture may reduce uncertainty for attackers and inadvertently improve adversarial planning. At the same time, opacity can preserve uncertainty and deterrence. The tradeoff is therefore not about “more versus less” information, but about how information is structured and how beliefs are formed.
Empirical evidence shows that defensive deception can remain effective even when attackers are explicitly informed that deception may be present. In the Tularosa Study, Ferguson-Walter et al. show that the presence of decoys, combined with disclosure that deception might be deployed, significantly impeded attacker progress, increased wasted effort, and altered adversarial behavior Ferguson-Walter et al. (2018). Participants who were aware of possible deception still triggered decoys at high rates and exhibited measurable decision disruption. These findings suggest that overt acknowledgment of deception does not eliminate its defensive value; rather, it can reshape attacker beliefs and incentives in ways that enhance deterrence and operational friction Rogers et al. (2026).
Recent results on the price of transparency clarify this distinction. Li and Zhu compare overt persuasion, where the signaling mechanism is publicly known, with covert signaling, where the mechanism is hidden Li and Zhu (2023). They show that transparency does not reduce the sender’s equilibrium payoff relative to opacity in a perfect Bayesian equilibrium. In strictly competitive settings, including zero-sum environments, transparent and opaque structures yield the same payoff. In other classes of games, however, opacity can introduce informational friction that degrades performance, and the gap between the two regimes can become significant. Thus, contrary to intuition, transparency can be strategically advantageous rather than costly.
A similar paradox appears in multi-agent systems. Additional feedback information can change equilibrium strategies in ways that do not necessarily reduce aggregate cost. As shown in differential games with open-loop versus closed-loop structures, the efficiency comparison depends on population size and cost parameters Başar and Zhu (2011). Effective communication design, therefore, requires understanding how information alters equilibrium incentives, rather than assuming that more feedback automatically improves coordination.
@@
Braess’ paradox illustrates that adding capacity to a network does not always improve overall performance Nagurney and Nagurney (2021); Di et al. (2014). When users act strategically, structural improvements can alter equilibrium behavior in ways that increase congestion rather than reduce it. In some cases, removing a link can improve system-wide outcomes.
A related observation appears in recent work on transportation resilience. The PRADA framework shows that informational perturbations, even when introduced with adversarial intent, can sometimes lead to improved equilibrium traffic conditions Yang et al. (2025c). These findings reinforce a broader lesson: system performance depends not only on infrastructure or controls, but on how strategic agents adapt to them.
In cyber systems, the same principle applies. Adding defensive layers or controls may unintentionally introduce complexity, shift attacker incentives, or create new bottlenecks. Resilience is therefore not a monotonic function of added resources. It emerges from the interaction between structure and strategic behavior.
@@
Surprise is central to cyber warfare, shaping strategy and amplifying operational impact. It exploits adversarial blind spots and introduces unexpected disruptions that undermine defenses. Unlike conventional warfare, where surprise is achieved through maneuvers or speed, cyber operations leverage stealth, misinformation, and novel attack vectors. APTs exemplify this, infiltrating quietly and striking only when conditions maximize the effect. Zero-day exploits deliver shock similarly by exploiting unknown vulnerabilities before defenses exist.
@@
Game-theoretic approaches formalize this interplay of deception and surprise. Cyber deception games, for instance, model how defenders mislead attackers about intentions and capabilities. Motivated by experiments on cognitive bias in red-team cyber range operations Beltz et al. (2026), in Yang and Zhu (2025a), cognitive triggers are designed to create microdeceptions that exploit bounded rationality, inducing adversaries into suboptimal choices by shaping their beliefs and expectations. This slows the attack process, forces attackers to waste additional effort, and ultimately creates strategic advantages for the defender.
Figure 1: A multi-echelon cyber warfare framework depicting the hierarchical structure of cyber conflict across policy, strategic, operational, and tactical levels. The cyber policy level defines long-term objectives through infrastructure development, alliances, and legal-normative frameworks such as the Tallinn Manual. These directives guide strategic planning, where campaigns progress through stages including reconnaissance, disruption, escalation, sustained assault, and strategic messaging. At the operational level, attacker and defender strategies evolve dynamically across stages indexed by $k$. Tactical actions (e.g., $T_{1}\rightarrow T_{2}\rightarrow T_{3}$) represent concrete maneuvers whose aggregated effects determine operational and strategic outcomes. The framework emphasizes cross-echelon interdependence: higher-level decisions shaping lower-level actions and realized outcomes feeding back to adjust strategy and policy, forming a coherent hierarchical meta-game.
@@ 4 Multi-Echelon Game-Theoretic Paradigm
Game theory provides the mathematical rigor to represent this hierarchy. Each echelon, policy, strategic, operational, tactical, and technical, corresponds to a different game with its own players, actions, payoffs, and equilibrium concepts. Yet, these games are not isolated: Equilibria at one level act as constraints or objectives for the next, while outcomes at lower levels feed back to reshape higher-level decisions. This recursive dependence gives rise to what we term a meta-game, in which stability requires consistency across all levels Ge and Zhu (2024); Zhu and Başar (2015).
@@ 4.1 Multi-Echelon Games
At the top of the hierarchy are cyber policy decisions. These include international agreements (e.g., cyber alliances, norms of behavior, and frameworks like the Tallinn Manual) and investments in national infrastructure. Such interactions can be modeled with coalition games Aumann and Dreze (1974). In this setting, the players are countries, $N=\{1,\dots,n\}$. Any subgroup of countries $S\subseteq N$ can form a coalition, and the coalition value function $v(S)$ describes the benefit of cooperation (e.g., shared defense, intelligence sharing). Game-theoretic solution concepts such as the core or the Shapley value tell us how to fairly divide the benefits of cooperation so that alliances remain stable.
@@
At the strategic level, cyber campaigns unfold over multiple stages, indexed by $k\in\{k_{1},k_{2},\dots,K\}$, as shown in Figure 1. Each stage corresponds to a different phase of operations: reconnaissance ($k=k_{1}$), disruption ($k=k_{2}$), escalation ($k=k_{3}$), sustained assault ($k=K$), and finally strategic messaging.
@@
As shown in the figure, the sequence of actions over time determines whether escalation occurs at stage $k=k_{3}$. Here, equilibrium concepts such as the $\epsilon$-Perfect Bayesian Nash Equilibrium Huang et al. (2020) are important: they model how attackers form beliefs when they are uncertain about which deception mechanism $\lambda$ is active.
@@
Finally, at the technical level, the focus is on developing new technological and computational solutions that directly enhance cyber defense. Large Language Models (LLMs) and agentic AI serve as key enablers, enabling adaptive responses, automated decision support, and the synthesis of vast streams of threat intelligence. For instance, LLM-driven agents can monitor logs, correlate signals across domains, and propose countermeasures in real time, while multi-agent AI frameworks enable distributed coordination among defensive components Zhu (2025a). In addition, adversarial learning techniques, such as Generative Adversarial Networks (GANs), support the creation of realistic honeypots and synthetic data sets that improve detection and resilience Gabrys et al. (2024). At this level, the emphasis is on computational innovation that equips defenders with tools to anticipate, deceive, and outmaneuver adversaries in highly dynamic environments.
@@ 4.2 Cross-Echelon Relationships
A canonical example of this setup is the Colonel Blotto game Roberson (2006); Hart (2008), in which two players simultaneously distribute forces across multiple battlefields, and outcomes at each battlefield are determined by relative strength. Here, the “battlefields” correspond to operational categories $\Lambda$, and the policy-derived weights $w^{*}$ adjust the relative importance of each category. Equilibrium allocations determine which operations dominate, are neutralized, or remain contested.
@@
At the operational level, a campaign is broken into phases $k=1,\dots,K$, each with a state $s^{k}$. The attacker chooses actions $a_{a}^{k}$ with strategy $\pi_{a}^{k}(s^{k})$, while the defender chooses actions $a_{d}^{k}$ with strategy $\pi_{d}^{k}(s^{k},\lambda)$, where $\lambda$ indexes which deception mechanism (e.g., honeypot, decoy) is deployed Gabrys et al. (2024); Huang and Zhu (2019a, 2020); Yang and Zhu (2025b). The defender’s cumulative payoff is
@@
At the technical level, $\theta$ captures the state of enabling technologies (e.g., advanced LLMs Touvron and others (2023); Team and DeepMind (2023) or GAN-based deception systems Goodfellow et al. (2014); Tolosana et al. (2020); Zellers et al. (2019)). Improvements at this level reduce costs $c_{d}(\theta)$, expand feasible action sets $A_{d}(\theta)$, alter the equilibrium strategies $(\xi_{d}^{*}(\theta),\xi_{a}^{*}(\theta))$ and their payoffs $U_{d}^{T},U_{a}^{T}$, and change the transition dynamics $T(\cdot;\theta)$. The associated equilibrium is denoted $E_{\text{technical}}=\theta^{*},$ representing the dominant technological state that parameterizes the equilibria at all higher levels. In this way, advances at the technical level propagate upward, shifting equilibria across every higher echelon of the hierarchy.